OpenStudy (anonymous):

Simplify: 16 – 4 + 14+ 2

OpenStudy (anonymous):

28

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

16-4=12 12+14=26 26+2=28

OpenStudy (anonymous):

i think 28

OpenStudy (anonymous):

thats not an option haha

OpenStudy (anonymous):

13 19 1 20

OpenStudy (anonymous):

what is the options then

OpenStudy (anonymous):

13 19 1 20

OpenStudy (anonymous):

Either your question is wrong or your answers are wrong.

OpenStudy (anonymous):

my life is wrong.

OpenStudy (anonymous):

check the question again.

OpenStudy (anonymous):

its not 13.... all it says is simplify w.e. screw it

OpenStudy (anonymous):

it's 28 hunn

OpenStudy (anonymous):

okay you guys are prob right. just not a stupid option

OpenStudy (anonymous):

did you mean Simplify: 16–4+14 /2?